# TIAM - A Metric for Evaluating Alignment in Text-to-Image Generation

### v1

@@ Acknowledgement
This work was granted access to the HPC resources of IDRIS under the allocation 2022-AD011014009 made by GENCI. This publication was also made possible by the use of the FactoryIA supercomputer, financially supported by the Ile-De-France Regional Council.
@@ Appendix A Proof of Proposition 1 and its generalization
We give the proof of proposition 1, that is easy to establish, then we derive a formula giving the number of templates in a more general realistic case (proposition 3). For this, we introduce an intermediate step (proposition 2) and most importantly a set of notation that facilitates the derivation of the final analytic formula.
Let us consider a collection of object sets $\{\mathcal{O}_{i}\}_{i=1}^{N}$ and a collection of attribute sets $\{\mathcal{A}_{i}\}_{i=1}^{N}$ that are used to define a template with $N$ objects, each being qualified by an attribute. Hence, at position $i$ (without considering the context), the prompts results from the inference of the template and contains a named object $o_{i}\in\mathcal{O}_{i}$ that is qualified by a (color) adjective $a_{i}\in\mathcal{A}_{i}$.
@@
Without attribute, each prompt contains $N$ objects that should be different thus the number of unique possible prompt is the number of (arrangements) $N$-permutations of $|\mathcal{O}|$ thus $\frac{|\mathcal{O}|!}{\left(|\mathcal{O}|-N\right)!}$. Similarly if one considers the attributes only that must be different at each position $i$ we have $N$-permutations of $|\mathcal{A}|$. Finally, since both are independent the final number of unique prompts generated by the template is the product of both. ∎
We can consider a generalization where the sets are different at each position, that is we remove the conditions $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$ from proposition 1. For example, such template could be defined to have a vehicle at first position, a fruit at the second position and an animal at the third one. The intersection of the sets at each position may also be non-empty, for example the vehicle may be $\{blue,red,green,black,yellow,white\}$, the fruit $\{red,green,yellow\}$ and the animal $\{green,black,yellow,white\}$. Similarly, the attributes can be repeated in the prompt e.g if one wants “a yellow car and a red apple and a red elephant”.
@@
if $\forall i\in[\![1,N]\!]$, $|\mathcal{O}_{i}|\geq N$, $|\mathcal{A}_{i}|\geq N$ thus the number of unique prompts generated by the template is $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. Hence, if $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$ thus the number of unique prompts generated by the template is $(|\mathcal{O}|.|\mathcal{A}|)^{N}$
@@
At any position $i$ any all the attributes $a_{i}\in\mathcal{A}_{i}$ can be associated with the object $o_{i}\in\mathcal{O}_{i}$ thus it gives $|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ possibilities. Since the repetition are allowed, we have a structure of arborescence (rooted tree), thus the total number of prompts is the product at any position, thus $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. ∎
@@
Table 5: Illustration of the set $\mathcal{P}$ (fourth col.) for the example of Table 4. The notation $\mathcal{U}_{i}^{0}$ means that any colored object of $\mathcal{U}_{i}^{0}$ can be used at that position
@@
Symmetrically, these configurations are also defined by the sets of positions where it appears only object at a unique position (thus from $\mathcal{U}_{i}^{0}$). We note $\mathcal{P}$ the set of these last positions (See Tab. 5 for an illustration of these sets). Note that the same element several times can appear several times in $\mathcal{P}$, actually when several objects with attributes can appear at the same position (e.g both “blue car” and “green apple” can appear at position 3 in our example). Formally, $\mathcal{P}$ is a set that contains the sets of indexes of the non-empty positions of all the N-uples that verify $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cap(\mathcal{U}_{j}\text{\textbackslash}\mathcal{U}_{j}^{0})\neq\emptyset$. Using this formalization of the problem, we have:
@@
For any configuration in $\mathcal{P}$, an object with attributes (from one of the $\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0}$) appear at most once since if $a_{i}=a_{j}$ thus $o_{i}\neq o_{j}$.
@@
Hence the exact number of prompts generated depends on the possible overlaps between objects with attributes, both the number of element and the place the can appear or not. By introducing more notations on these elements and their position, one could derive a formula but it would be tedious, without obvious interest in practice. Indeed, $\mathcal{P}$ can be built easily through a tree structure, by considering iteratively all the sets $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cup\{\emptyset\}$, and allowing only the nodes that did not previously appear and the branch corresponding to non-empty set in $\mathcal{P}$ (such as in Tab. 5).
@@ B.1 Without attribute
The 24 COCO labels used for the general study of sections 4.1 and 4.2 are the following: bicycle, car, motorcycle, truck, fire hydrant, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, banana, apple, broccoli, carrot, chair, couch, oven, refrigerator. The prompt template is “a photo of $det(o_{1})$ $o_{1}$ and $det(o_{2})$ $o_{2}$”.
@@
We do not use commas because it tends to reduce the score. We report in Tab 6 the comparison between the prompt “a photo of $det(o_{1})$ $o_{1}$ next to $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$” and the prompt with a comma “a photo of $det(o_{1})$ $o_{1}$, $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$”.
@@ B.2 With Attribute
Appendix C Reference colors
We plot on Fig 13 the interpolation of the best examples in Lab on the CIE 1931 Chromaticity Diagram.
@@ Appendix C Reference colors
We do not use our reference colors orange and brown because they are respectively too close to red and black in the CIELab space (Fig. 12).
@@ Appendix D Occurrence of objects on images
We show in our experiments (section 4.3) that the initial objects in the template tend to appear more frequently than objects inserted subsequently and reinforce the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image.
We present the result for two objects in Fig. 14 and for three objects in Fig. 15.
Figure 14: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt includes two objects.
Figure 15: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt includes three objects.
@@ Appendix G Text-to-image models Setup
It exists a different configuration of the Deepfloyd IF. We use for the first stage the L version [^11] https://huggingface.co/DeepFloyd/IF-I-L-v1.0 with 100 inference steps and the second stage the M version [^12] https://huggingface.co/DeepFloyd/IF-II-M-v1.0 with 50 inference steps. We use both the DDPMS scheduler and guidance scale of 7 for the first stage and 4 for the second stage. With this configuration, we produce images of size 256$\times$256.
@@ Appendix H Semantic Link
For all the models we observe some clusters of objects from the same macro-class, in particular animals. It shows that when two objects are semantically close, they tend to be harder to be generated in the same image. Tang et al. [27] obtain results in the same vein, showing that it is easier to generate a non-cohyponym than a cohyponym.
However, the effect remains slight, suggesting that the cohyponymy has either an indirect or minor link to this difficulty of generation. We quantified the effect by computing the correlation between the TIAM score for all the templates with two objects and the semantic distance between the two objects. We used various methods to estimate the semantic distance, including Wu-Palmer, the CLIPscore, and the cosine similarity between the embedding of the token in the prompt (before the attention of the transformer) for SD 1.4 and SD 2. We report the results in Tab 7. For all distances, these correlations are negative (confirming the effect) but their absolute values are less than 0.3 (confirming the effect is slight).
Correlation with SD 1.4 SD 2
CLIPscore -0.09 -0.1
Embedding matrix -0.06 -0.12
Wu-Palmer -0.28 -0.28
Similarity in context -0.17 -0.24
Key (mean±$\pm\sigma$) $-0.17\pm{0.08}$ $-0.17\pm{0.09}$
Value (mean±$\pm\sigma$) $-0.19\pm{0.07}$ $-0.17\pm{0.08}$
Table 7: Correlation between the semantic distance between two objects ($o_{1}$ and $o_{2}$) and the alignment of the image generated by Stable Diffusion TIAM and a template including $o_{1}$ and $o_{2}$. The objects belong to three macro-classes (vehicles, animals, foods). The CLIPscore is the cosine similarity between the CLIP textual features (first eot embedding) of $o_{1}$ and $o_{2}$. Embedding matrix is the cosine similarity between the representation of tokens $o_{1}$ and $o_{2}$ in the semantic space prior to attention in the CLIP model. Wu-palmer stands for the Wu-Palmer similarity between $o_{1}$ and $o_{2}$. The similarity in context is the cosine similarity between the representation of the tokens $o_{1}$ and $o_{2}$ in the last hidden states of the embedding of the prompt by the CLIP textual encoder. Finally, Key and Value indicate that we compute the average score and standard deviation ($\sigma$) of all the correlations of the representations Key and Value. We calculated these correlations for each attention head of each layer of the U-net.
@@ Appendix I Attribute binding
We report the TIAM per object on Fig. 22, showing that the first object is more often generated and correctly colored. We compute the binding success rate, but by differentiating by colors for attributes in the first position (Fig. 23) and attributes in the second position (Fig. 24). We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
@@
We remind the architecture of the Latent Diffusion Model of Rombach et al. (2021) in Figure 26, since the theoretical explanation of Section 4.2 of the main paper relies on it.
Figure 25: Standardized score per seed for SD 1.4 and SD 2. We observe that both of them have globally the same “good” and “bad” seeds (see subsection 4.2 for explanations).
@@
During the investigation of the seed performances of the models, we made a noteworthy finding. We observed that the SD models exhibited similar behavior in relation to the seeds (if we do not consider the performance gap between the score of the model) i.e. when we standardize the score of each seed for each model (Fig 25) the score exhibits a remarkable similar trend for both models with the same “good” and “bad” seeds. However, we explain in the article (section 4.2) that the “good” and “bad” seeds are specific to each model, which may seem contradictory at first glance.
@@
To explain this point, we need to remind the training of diffusion models. Let $x_{0}$ be an original image. We had on the image a scaled quantity of noise $\epsilon\sim\mathcal{N}(0,1)$ to obtain $x_{t}$ ($\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$) and the U-net $\epsilon_{\theta}$ try to predict with the added noise (the loss function ($\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$). At each training step a $t$ is drawn the model must predict the noise. In the case of LDM, just replace the $x$ with $z$ because the diffusion process is in the latent space.
@@
As the two models are trained on the same data, we hypothesize that they are trained on the same $z_{t}$ i.e. they are trained to predict the same $\epsilon$. However $\epsilon$ is random, we suspect that all the $z_{t}$ are precomputed and the associate $\epsilon$ saved, and they used the same for the training of both models. Because of this particularity during the training, the models try to retrieve the $z_{0}$ from multiple possible $t$. The models learn a similar path of reverse diffusion. At inference, when we draw a similar noise $\chi_{T}$, the reverse process will be similar and conduct to near $\chi_{t}$ and finally to a near $\chi_{0}$. We show in Fig.8 that with the same prompt and the same starting noise we exhibit strong composition similarity that explains the parallel performance of the seed.
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]

### v2

@@ Acknowledgement
This work was granted access to the HPC resources of IDRIS under the allocation 2022-AD011014009 made by GENCI. This publication relied on the use of the FactoryIA supercomputer, financially supported by the Ile-de-France Regional Council.
@@ Appendix A Proof of Proposition 1 and its Generalization
We give the proof of proposition 1, which is easy to establish, then we derive a formula giving the number of templates in a more general realistic case (proposition 3). For this, we introduce an intermediate step (proposition 2) and most importantly a set of notations that facilitates the derivation of the final analytic formula.
Let us consider a collection of object sets $\{\mathcal{O}_{i}\}_{i=1}^{N}$ and a collection of attribute sets $\{\mathcal{A}_{i}\}_{i=1}^{N}$ that are used to define a template with $N$ objects, each being qualified by an attribute. Hence, at position $i$ (without considering the context), the prompt results from the inference of the template and contains a named object $o_{i}\in\mathcal{O}_{i}$ that is qualified by a (color) adjective $a_{i}\in\mathcal{A}_{i}$.
@@
Without attribute, each prompt contains $N$ objects that should be different; thus the number of unique possible prompts is the number of (arrangements) $N$-permutations of $|\mathcal{O}|$ thus $\frac{|\mathcal{O}|!}{\left(|\mathcal{O}|-N\right)!}$. Similarly, if one considers the attributes only that must be different at each position $i$, we have $N$-permutations of $|\mathcal{A}|$. Finally, since both are independent, the final number of unique prompts generated by the template is the product of both. ∎
We can consider a generalization where the sets are different at each position, that is we remove the conditions $a_{i}\neq a_{j}$ and $o_{i}\neq o_{j}$ from proposition 1. For example, such a template could be defined to have a vehicle at the first position, a fruit at the second position, and an animal at the third one. The intersection of the sets at each position may also be non-empty, for example, the vehicle may be $\{blue,red,green,black,yellow,white\}$, the fruit $\{red,green,yellow\}$, and the animal $\{green,black,yellow,white\}$. Similarly, the attributes can be repeated in the prompt e.g. if one wants “a yellow car and a red apple and a red elephant”.
@@
if $\forall i\in[\![1,N]\!]$, $|\mathcal{O}_{i}|\geq N$, $|\mathcal{A}_{i}|\geq N$ thus the number of unique prompts generated by the template is $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. Hence, if $\forall i\in[\![1,N]\!],\mathcal{A}_{i}=\mathcal{A},\mathcal{O}_{i}=\mathcal{O}$, thus the number of unique prompts generated by the template is $(|\mathcal{O}|.|\mathcal{A}|)^{N}$
@@
At any position $i$ any all the attributes $a_{i}\in\mathcal{A}_{i}$ can be associated with the object $o_{i}\in\mathcal{O}_{i}$ thus it gives $|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$ possibilities. Since repetitions are allowed, we have a structure of arborescence (rooted tree), thus the total number of prompts is the product at any position, thus $\prod_{i=1}^{N}|\mathcal{A}_{i}|.|\mathcal{O}_{i}|$. ∎
@@
Table 5: Illustration of the set $\mathcal{P}$ (fourth col.) for the example of Tab. 4. The notation $\mathcal{U}_{i}^{0}$ means that any colored object of $\mathcal{U}_{i}^{0}$ can be used at that position
@@
Symmetrically, these configurations are also defined by the sets of positions where only one object appears at a single position (thus from $\mathcal{U}_{i}^{0}$). We note $\mathcal{P}$ the set of these last positions (See Tab. 5 for an illustration of these sets). Note that the same element can appear several times in $\mathcal{P}$, actually when several objects with attributes can appear at the same position (e.g. both “blue car” and “green apple” can appear at position 3 in our example). Formally, $\mathcal{P}$ is a set that contains the sets of indexes of the non-empty positions of all the N-uples that verify $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cap(\mathcal{U}_{j}\text{\textbackslash}\mathcal{U}_{j}^{0})\neq\emptyset$. Using this formalization of the problem, we have:
@@
For any configuration in $\mathcal{P}$, an object with attributes (from one of the $\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0}$) appears at most once since if $a_{i}=a_{j}$ thus $o_{i}\neq o_{j}$.
@@
Hence the exact number of prompts generated depends on the possible overlaps between objects with attributes, both the number of elements and the place they can appear or not. By introducing more notations on these elements and their position, one could derive a formula but it would be tedious, without obvious interest in practice. Indeed, $\mathcal{P}$ can be built easily through a tree structure, by considering iteratively all the sets $(\mathcal{U}_{i}\text{\textbackslash}\mathcal{U}_{i}^{0})\cup\{\emptyset\}$, and allowing only the nodes that did not previously appear and the branch corresponding to a non-empty set in $\mathcal{P}$ (such as in Tab. 5).
@@ B.1 Without attribute
The 24 COCO labels used for the study of Sections 4.1 and 4.2 are the following: bicycle, car, motorcycle, truck, fire hydrant, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, banana, apple, broccoli, carrot, chair, couch, oven, refrigerator. The prompt template is “a photo of $det(o_{1})$ $o_{1}$ and $det(o_{2})$ $o_{2}$”.
@@
We do not use commas because it tends to reduce the score. We report in Tab. 6 the comparison between the prompt “a photo of $det(o_{1})$ $o_{1}$ next to $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$” and the prompt with a comma “a photo of $det(o_{1})$ $o_{1}$, $det(o_{2})$ $o_{2}$ and $det(o_{3})$ $o_{3}$”.
@@ B.2 With Attribute
Appendix C Reference Colors and Other Possible Attributes (Size, Texture)
We plot in Fig. 13 the interpolation of the best examples in Lab on the CIE 1931 Chromaticity Diagram. The exact chroma values were extracted from Fig.1 in [25], available at this link.
@@ Appendix C Reference Colors and Other Possible Attributes (Size, Texture)
We do not use our reference colors orange and brown because they are too close to red and black respectively in the CIELab space (Fig. 12).
@@
To consider an attribute in TIAM, two crucial points need to be considered:
  • being able to extract the attribute from the image, with a sufficient level of reliability
  • being able to name the attributes with unambiguous words in the prompt, ideally in several languages
The attribute colors have the advantage of quite easily meeting these two conditions, as explained in Section 3.2 of the main paper, in particular, thanks to the works of Berlin and Kay [3]. In other cases, however, this may prove trickier.
To extract the attribute size, one can rely on the bounding box of the object detector. However, to determine whether an object is large, medium, or small for example, it may also require to estimate its depth in the image (that can be done from monocular images to a certain extent [43]), possibly an estimation of the intrinsic parameter matrix of the camera (while the image is generated) as well as the knowledge of the typical dimensions of the considered object or living being. Each of these estimations is a potential source of approximation that challenges the reliability of the final judgment.
Naming the size may also be diverse. In the example above, large may be replaced by big in English and small by tiny in some cases. One could imagine relying on lists of synonyms or setting a threshold on the similarity to the textual embedding (e.g. BERT) of an arbitrary predefined list of possible sizes (e.g. large, medium, small), but there is no guarantee to get a list as unambiguous as in the case of colors. To our knowledge, there is no equivalent of the study of Berlin and Kay in that case. Moreover, references to sizes tend to include intensifiers more frequently than references to colors, like in expressions such as very small or fairly large, which adds an element of diversity. Finally, the way sizes are expressed can depend on forms of collocations. For instance, we can refer to a “tall man” but not to a “tall balloon”. While this example can be probably explained by the form of the object, the easiest way to deal with such problem would be to collect co-occurrences from a corpus for (size adjective, noun) pairs and to use these co-occurrences as a filter after the generation of a prompt.
Extracting texture is a long-term and well-known task in computer vision [24] and many methods have been proposed to address it [22]. The question of naming the texture with unambiguous names may seem delicate. In the famous Brodatz dataset for example (available here), one can note that several of them are named Wood shingle roof (here and here), Brick wall (here, here, or here) or Sand (here, here and here) among others. Bhushan et al. (1997) nevertheless identified a list of 98 representative words used to describe texture in English [4] that was further reduced to 47 in the Describable Textures Dataset [8]. However, such a number remains quite large (much more than the 11 colors of Berlin and Kay), such that the correspondence in other language than English is hazardous, not to mention the fact that some of them may seem ambiguous to several human users (our human study in Section K includes users less than 10 years old with a limited vocabulary, as well as a majority of persons that are not computer vision scientists). Using Multidimensional Scaling (MDS) on these 98 words, Bhushan et al. (1997) nevertheless identified 11 clusters that could be used, although naming these clusters is still problematic in practice. Finally, the easier approach may be to use three axes, identified with the MDS as well, namely:
  • repetitive versus nonrepetitive textures
  • the nature of orientation: linearly oriented textures $\to$ multiple or no orientation $\to$ circularly oriented textures.
  • complexity or simplicity of the surface
Hence, considering other attributes than colors in TIAM is likely feasible, but integrating them neatly into the metric may require some work.
@@ Appendix D Occurrence of Objects on Images
We show in our experiments (Section 4.3) that the initial objects in the template tend to appear more frequently than objects inserted subsequently and reinforce the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image. We present the result for two objects in Fig. 14 and three objects in Fig. 15.
Figure 14: The proportion of occurrences of each object, based on its position in the prompt. The template of the prompt includes two objects.
Figure 15: The proportion of occurrences of each object, based on its position in the prompt. The template of the prompt includes three objects.
@@ Appendix G Text-to-Image Models Setup
It exists a different configuration of the Deepfloyd IF. We use for the first stage the L version [^8] https://huggingface.co/DeepFloyd/IF-I-L-v1.0 with 100 inference steps and for the second stage the M version [^9] https://huggingface.co/DeepFloyd/IF-II-M-v1.0 with 50 inference steps. We use both the DDPMS scheduler and guidance scale of 7 for the first stage and 4 for the second stage. With this configuration, we produce images of size 256$\times$256.
@@ Appendix H Semantic Link
For all the models, we observe some clusters of objects from the same macro-class, in particular animals. It shows that when two objects are semantically close, they tend to be harder being generated in the same image. Tang et al. [44] obtain results in the same vein, showing that it is easier to generate a non-cohyponym than a cohyponym.
However, the effect remains slight, suggesting that the cohyponymy has either an indirect or minor link to this difficulty of generation. We quantified the effect by computing the correlation between the TIAM score for all the templates with two objects and the semantic distance between the two objects. We used various methods to estimate the semantic distance, including Wu-Palmer, the CLIPscore, and the cosine similarity between the embedding of the token in the prompt (before the attention of the transformer) for SD 1.4 and SD 2. For all distances, these correlations were negative (confirming the effect) but their absolute values were less than 0.5 (confirming the effect is slight).
@@ Appendix I Attribute Binding
We report the TIAM per object in Fig. 22, showing that the first object is more often generated and correctly colored. We compute the binding success rate, but by differentiating by colors for attributes in the first position (Fig. 23) and attributes in the second position (Fig. 24). We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
@@
We remind the architecture of the Latent Diffusion Model of Rombach et al. (2021) in Fig. 25, since the theoretical explanation of Section 4.2 of the main paper relies on it.
@@
During the investigation of the seed performances of the models, we made a noteworthy finding. We observed that the SD models exhibited similar behavior in relation to the seeds (if we do not consider the performance gap between the score of the model) i.e. when we standardize the score of each seed for each model (Fig. 26) the score exhibits a remarkably similar trend for both models with the same “good” and “bad” seeds. However, we explain in the article (Section 4.2) that the “good” and “bad” seeds are specific to each model, which may seem contradictory at first glance.
Figure 26: Standardized score per seed for SD 1.4 and SD 2. We observe that both of them have globally the same “good” and “bad” seeds (see Section J for explanations).
@@
To explain this point, we need to remember the training of diffusion models. Let $x_{0}$ be an original image. We had on the image a scaled quantity of noise $\epsilon\sim\mathcal{N}(0,1)$ to obtain $x_{t}$ ($\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$) and the U-net $\epsilon_{\theta}$ try to predict with the added noise (the loss function ($\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$). At each training step a $t$ is drawn the model must predict the noise. In the case of LDM, just replace the $x$ with $z$ because the diffusion process is in the latent space.
@@
As the two models are trained on the same data, we hypothesize that they are trained on the same $z_{t}$ i.e. they are trained to predict the same $\epsilon$. However $\epsilon$ is random, we suspect that all the $z_{t}$ are precomputed and the associate $\epsilon$ saved, and they used the same for the training of both models. Because of this particularity during the training, the models try to retrieve the $z_{0}$ from multiple possible $t$. The models learn a similar path of reverse diffusion. At inference, when we draw a similar noise $\chi_{T}$, the reverse process will be similar and conduct to near $\chi_{t}$ and finally to a near $\chi_{0}$. We show in Fig. 27 that with the same prompt and the same starting noise, we exhibit strong composition similarity that explains the parallel performance of the seed.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption